# A CMC existence result for expanding cosmological spacetimes

### v2

@@ Abstract
We establish a new CMC (constant mean curvature) existence result for cosmological spacetimes, i.e., globally hyperbolic spacetimes with compact Cauchy surfaces satisfying the strong energy condition. If the spacetime contains an expanding Cauchy surface and is future timelike geodesically complete, then the spacetime contains a CMC Cauchy surface. This result settles, under certain circumstances, a conjecture of the authors and a conjecture of Dilts and Holst. Our proof relies on the construction of barriers in the support sense, and the CMC Cauchy surface is found as the asymptotic limit of mean curvature flow. Analogous results are also obtained in the case of a positive cosmological constant $\Lambda>0$. Lastly, we include some comments concerning the future causal boundary for cosmological spacetimes which pertain to the CMC conjecture of the authors.
@@ 1 Introduction
Let $(M,g)$ be a spacetime with compact Cauchy surfaces, and suppose there exists a smooth spacelike Cauchy surface with mean curvature $H\geq 0$. If $(M,g)$ is future timelike geodesically complete and satisfies the strong energy condition, i.e., ${\rm Ric}(X,X)\geq 0$ for all timelike $X$, then $(M,g)$ contains a CMC Cauchy surface.
@@ 2 Spacetime and mean curvature flow preliminaries
We recall some basic definitions and facts. A spacetime is a smooth time-oriented connected Lorentzian manifold $(M,g)$. Classically, a spacetime $(M,g)$ is globally hyperbolic if it is strongly causal and all ‘causal diamonds’ $J^{+}(p)\cap J^{-}(q)$ are compact. A Cauchy (hyper)surface for a spacetime $(M,g)$ is a set $S$ that is met by every inextendible causal curve exactly once, which implies that $S$ is a topological (in fact locally Lipschitz) hypersurface and $M\approx\mathbb{R}\times S$ topologically. It’s a well-known fact that a spacetime $(M,g)$ is globally hyperbolic if and only if it contains a Cauchy surface $S$.
Fix a globally hyperbolic spacetime $(M,g)$. Let $F_{0}\colon V\to M$ be an embedding such that $V_{0}:=F_{0}(V)$ is a smooth compact spacelike Cauchy surface. Consider the evolution problem
@@
and $H(x,s)$ is the mean curvature of $V_{s}$ at $F(x,s)$ defined with respect to $\nu(x,s)$, and $c\in\mathbb{R}$ is a constant. By [11, Thm. 4.1], a solution to this evolution problem exists at least locally for $s\in[0,s_{0})$; that is, there is a smooth map
@@
Moreover, if the flow $\{V_{s}\}_{s\in[0,s_{0})}$ remains in a compact region of the spacetime $M$, then the solution can be extended beyond $s_{0}$. [^3] More generally, [11, Thm. 4.1] allows not just for a constant $c$ in (2.4), but for any smooth function $\mathscr{H}$ on $M$ satisfying a certain monotonicity condition.
@@
which follows from (2.1) and (2.7) (cf. [11, Eq. 5]).
@@ 3 Proof of Theorem 3
Recall that for each $q\in J^{+}(S)$, the Lorentzian distance function $d(S,q)$ is defined as the supremum of $L(\gamma)$ ($=$ length of $\gamma$) over all causal curves $\gamma$ from $S$ to $q$.
@@
That $S_{\tau}$ is a Cauchy surface follows from results in [17]; see especially Lemmas 3.12 and 3.15 in that reference.
@@
We show $S_{\tau}$ has mean curvature $H\leq\tfrac{n}{\tau}$ in the support sense. Since $S$ is compact and $(M,g)$ is globally hyperbolic, the function $x\mapsto d(S,x)$ is continuous, finite-valued, and for each $q\in S_{\tau}$, there is a unit speed timelike geodesic $\gamma\colon[0,\tau]\to M$ that realizes the distance from some $p\in S$ to $q$, i.e., $L(\gamma)=d(S,q)$. In particular, each segment of $\gamma$ is maximal and so there are no cut points to $\gamma(t)$ along $\gamma|_{[t,\tau]}$, for $t>0$. Consequently, each future Lorentzian sphere $\Sigma_{u}=\{x\in M\mid d\big{(}\gamma(t),x\big{)}=u\}$, for $u\in(0,\tau-t]$, is smooth near $\gamma(u+t)$. It follows that $\Sigma_{\tau-t}$ lies locally to the future of $S_{\tau}$ near $q$, for otherwise there would be a causal curve from $S$ to $S_{\tau}$ with length greater than $\tau$. Hence $\Sigma_{\tau-t}$ is a future support hypersurface for $S_{\tau}$ at $q$. Let $H(u)$ be the mean curvature of $\Sigma_{u}$ at $\gamma(u+t)$. The Raychaudhuri equation (for a hypersurface orthogonal geodesic vector field) gives, when restricted to $\gamma$,
$\frac{dH}{du}\leq-{\rm Ric}(\gamma^{\prime},\gamma^{\prime})-\frac{H^{2}}{n}\leq-\frac{H^{2}}{n}.$ (3.9)
@@
Standard ODE comparison results yield $H(u)\leq\frac{n}{u}$ (see e.g., [23, Section 1.6]), [12, Section 2]). Evaluating at $u=\tau-t$ and letting $t\to 0$ gives the result. ∎
@@
Remark. In fact for $\tau$ small, and provided that $S$ is smooth and spacelike, $S_{\tau}$ will be a Cauchy surface even without the completeness assumption. More concretely, define the “future existence time” $T_{0}>0$ from $S$ to be the largest value (possibly =absent=$\infty$= ∞) such that $\exp_{p}(tu)$ is defined for all $p\in S$ and all $t<T_{0}$. (Here $u$ is the future timelike unit normal field to $S$.) Then, by essentially the same argument, $S_{\tau}$ will be a Cauchy surface for all $\tau<T_{0}$. In this case, for such an $S$, the conclusion of Proposition 4 holds for all $\tau<T_{0}$.
@@
Let $(M,g)$ be a spacetime with compact Cauchy surfaces. Suppose that $S_{1}$ is a smooth and spacelike Cauchy surface with mean curvature $H_{1}>c$ and that $S_{2}\subset I^{+}(S_{1})$ is a (not necessarily smooth) Cauchy surface with mean curvature $H_{2}\leq b$ in the support sense, for some constant $b<c$. Then there exists a smooth spacelike Cauchy surface $S$ between [^5] I.e., $S\subset I^{+}(S_{1})\cap I^{-}(S_{2})$ $S_{1}$ and $S_{2}$ with constant mean curvature $H=c$, which is obtained as the asymptotic limit of mean curvature flow.
@@
Let $V_{0}=F_{0}(V)$ be any embedded smooth spacelike Cauchy surface between $S_{1}$ and $S_{2}$. Consider the evolution problem (2.1). We show that the corresponding flow $V_{s}$, given by (2.2), always lies within the compact region between $S_{1}$ and $S_{2}$. Indeed, suppose there is a point $(x_{0},s_{0})\in V\times[0,\infty)$ where the flow $V_{s}$ touches $S_{2}$ for the first time. At this point $p_{0}:=F(x_{0},s_{0})$, we must have $\tfrac{d}{ds}u\geq 0$, as otherwise the flow $V_{s}$ would necessarily have touched $S_{2}$ at some value $s_{1}<s_{0}$. We conclude from (2.8) that $H(x_{0},s_{0})\geq c$. Since $S_{2}$ has mean curvature $H_{2}\leq b<c$ in the support sense, for each $\varepsilon>0$, there exists a future support hypersurface $W$ for $S_{2}$ at $p_{0}$ with mean curvature $H_{W}\leq b+\varepsilon$. By choosing $\varepsilon$ sufficiently small, it’s easy to see that there is a $d\in(b,c)$ such that $H_{V_{s_{0}}}>d$ at $p_{0}$ and $H_{W}<d$ at $p_{0}$. Since $W$ is locally to the future of $S_{2}$ near $p_{0}$ and hence also locally to the future of $V_{s_{0}}$ near $p_{0}$, we arrive at a contradiction to the maximum principle. Thus the flow always remains below $S_{2}$. Likewise, the flow always remains above $S_{1}$.
@@
Since the Cauchy surfaces are compact, we can find $t$-slices of the temporal function 2.5 such that $S_{1}$ always lies above a slice $t=\alpha$ and $S_{2}$ lies below a slice $t=\beta$. Therefore the flow always remains in the smooth compact region of spacetime between $t=\alpha$ and $t=\beta$. It follows via [11, Thms. 4.1 and 4.2] that the flow $V_{s}$ is defined for all $s\in[0,\infty)$ and (by taking a suitable subsquence) limits to a smooth spacelike Cauchy surface $S$ with constant mean curvature $H=c$. ∎
@@
Remark. It’s clear from the above proof that Proposition 5 generalizes to the case when $S_{1}$ has mean curvature $H_{1}\geq a$ also in the support sense, for some $a>c$. An existence result similar to [11, Thm. 4.2] was obtained by Claus Gerhardt [19], again assuming smooth barriers. Gerhardt’s proof was via a variational method and did not require the Ricci curvature condition. There appears to be certain technical issues in trying to extend Gerhardt’s proof to the case of barriers in the support sense. A closely related situation arises in [1]. We thank the authors of [1] for a discussion pertaining to some of these issues.
@@
Proof of Theorem 3. By assumption, there is a smooth spacelike Cauchy surface $V$ with mean curvature $H_{V}\geq 0$. If $H_{V}=0$ identically, then we’re done. Otherwise, there is a point $p\in V$ with $H_{V}(p)>0$. We can perturb $V$ slightly to the past to produce a smooth spacelike Cauchy surface with strictly positive mean curvature. To see this, let $V_{t}$ be a variation of $V_{0}=V$ with variation vector $\phi\nu$, where $\phi$ is a smooth function on $V$ and $\nu$ is the future directed unit timelike normal to $V$. (Here we’re abusing notation and letting $t$ denote the flow parameter of the variation and not the temporal function given in (2.5).) Let $H_{t}$ be the mean curvature of $V_{t}$. A standard computation shows that
@@
$\frac{\partial H_{t}}{\partial t}\bigg{|}_{t=0}\,=\,-L\phi\quad\text{ where }\quad L\phi\,:=\,-\Delta\phi+\big{(}{\rm Ric}(\nu,\nu)+\tfrac{H_{0}^{2}}{n}+\sigma^{2}\big{)}\phi,$
@@
where $\Delta$ is the spatial Laplacian on $V$ and $\sigma^{2}$ is norm squared of the trace-free part of the second fundamental form of $V$. Since $H_{0}(p)>0$, the Rayleigh quotient formula for the principal eigenvalue $\lambda_{1}$ of $L$ shows that $\lambda_{1}>0$. Then, by varying with respect to the principal eigenfunction, there are times $t<0$ such that $H_{t}>0$ everywhere on $V_{t}$.
@@
Thus we can assume – without loss of generality – that $H_{V}>0$. Choose $c>0$ such that $H_{V}>c$. By taking $\tau$ sufficiently large, Proposition 4 implies the existence of a Cauchy surface $V_{\tau}$ with mean curvature $H_{\tau}<c$ in the support sense. Then Proposition 5 implies that there is a smooth spacelike Cauchy surface between $V$ and $V_{\tau}$ with constant mean curvature $H=c$. ∎

### v3

@@ Abstract
We establish a new CMC (constant mean curvature) existence result for cosmological spacetimes, by which we mean globally hyperbolic spacetimes with compact Cauchy surfaces. If a cosmological spacetime satisfying the strong energy condition contains an expanding Cauchy surface and is future timelike geodesically complete, then the spacetime contains a CMC Cauchy surface. This result settles, under certain circumstances, a conjecture of the authors and a conjecture of Dilts and Holst. Our proof relies on the construction of barriers in the support sense, and the CMC Cauchy surface is found as the asymptotic limit of mean curvature flow. Analogous results are also obtained in the case of a positive cosmological constant $\Lambda>0$. Lastly, we include some comments concerning the future causal boundary for cosmological spacetimes which pertain to the CMC conjecture of the authors.
@@ 1 Introduction
Let $(M,g)$ be a spacetime with compact Cauchy surfaces, and suppose there exists a smooth spacelike Cauchy surface $V$ with mean curvature $H\geq 0$. If $(M,g)$ is future timelike geodesically complete and satisfies the strong energy condition, i.e., ${\rm Ric}(X,X)\geq 0$ for all timelike $X$, then $(M,g)$ contains a CMC Cauchy surface.
@@ 2 Spacetime and mean curvature flow preliminaries
We recall some basic definitions and facts. By a spacetime, we mean an $(n+1)$-dimensional, $n\geq 2$, time-orientable Lorentzian manifold $(M,g)$, which, for the analytic purposes of this paper, we assume smoothness of both the manifold and metric. Classically, a spacetime $(M,g)$ is globally hyperbolic if it is strongly causal and all ‘causal diamonds’ $J^{+}(p)\cap J^{-}(q)$ are compact. A Cauchy (hyper)surface for a spacetime $(M,g)$ is a set $S$ that is met by every inextendible causal curve exactly once, which implies that $S$ is a topological (in fact locally Lipschitz) hypersurface and $M$ is homeomorphic to $\mathbb{R}\times S$. It is a well-known fact that a spacetime $(M,g)$ is globally hyperbolic if and only if it contains a Cauchy surface $S$.
Fix a globally hyperbolic spacetime $(M,g)$. Let $F_{0}\colon V\to M$ be an embedding such that $V_{0}:=F_{0}(V)$ is a smooth compact spacelike Cauchy surface. The following evolution problem defines mean curvature flow with prescribed constant mean curvature $c\in\mathbb{R}$:
@@
and $H(x,s)$ is the mean curvature of $V_{s}$ at $F(x,s)$ defined with respect to $\nu(x,s)$. By [12, Thm. 4.1], a solution to this evolution problem exists at least locally for $s\in[0,s_{0})$; that is, there is a smooth map
@@
Moreover, if the flow $\{V_{s}\}_{s\in[0,s_{0})}$ remains in a compact region of the spacetime $M$, then the solution can be extended beyond $s_{0}$. [^3] More generally, [12, Thm. 4.1] allows not just for a constant $c$ in (2.4), but for any smooth function $\mathscr{H}$ on $M$ satisfying a certain monotonicity condition.
@@
which follows from (2.1) and (2.7) (cf. [12, Eq. 5]).
@@ 3 Proof of Theorem 3
Recall that for each $q\in J^{+}(S)$, the Lorentzian distance function $d(S,q)$ is defined as the supremum of $L(\gamma)$ (== length of $\gamma$) over all causal curves $\gamma$ from $S$ to $q$.
@@
For the convenience of the reader, we provide a complete proof that $S_{\tau}$ is a Cauchy surface (cf. [18, Lem. 3.15]). We will use the following well-known fact: Since $S$ is compact and $(M,g)$ is globally hyperbolic, the function $x\mapsto d(S,x)$ is continuous, finite-valued, and for each $q\in S_{\tau}$, there is a unit speed timelike geodesic $\gamma\colon[0,\tau]\to M$ that realizes the distance from some $p\in S$ to $q$, i.e., $L(\gamma)=d(S,q)$.
$S_{\tau}$ is closed by continuity of the distance function. Moreover, $S_{\tau}$ is acausal: Clearly $S_{\tau}$ is achronal. Let $p,q,\gamma$ be as in the first paragraph of this proof. If there is another point $q^{\prime}\in S_{\tau}$ with $q^{\prime}\in J^{+}(q)$, then there is a null geodesic $\lambda$ from $q$ to $q^{\prime}$. Consider a convex neighborhood $U$ about $q$. Let $x,y\in U$ be points such that $x\in\gamma$ and $y\in\lambda$. The geodesic $\sigma\subset U$ from $x$ to $y$ has length strictly greater than the portion of $\gamma$ from $x$ to $q$ [27, Prop. 5.34]. Therefore the causal curve formed by following $p$ to $x$ along $\gamma$, then $x$ to $y$ along $\sigma$, and lastly $y$ to $q^{\prime}$ along $\lambda$ has length strictly greater than $\tau$ – a contradiction. Thus $S_{\tau}$ is closed and acausal. Moreover, it can be seen that $S_{\tau}$ has no edge points and consequently is a topological hypersurface [27, Cor. 14.26].
Now we show that $S_{\tau}$ is compact. Let $h$ be a complete Riemannian metric on $M$. If $S_{\tau}$ is not compact, then there is a sequence of points $q_{n}\in S_{\tau}$ such that the $h$-distance from $S$ to $q_{n}$ tends to infinity. Let $\gamma_{n}\colon[0,b_{n}]\to M$ be an $S$-maximizer from $p_{n}\in S$ to $q_{n}$, parameterized by $h$-arclength so that $b_{n}\to\infty$. Since $S$ is acausal, each $\gamma_{n}$ is a timelike geodesic (when appropriately reparameterized). Let $\widetilde{\gamma}_{n}\colon[0,\infty)\to M$ be the future inextendible timelike geodesic extension of $\gamma_{n}$, also parameterized by $h$-arclength. By compactness of $S$, there is a $p\in S$ such that $p_{n}\to p$ (after passing to a subsequence). The limit curve lemma guarantees the existence of a future inextendible causal curve $\gamma\colon[0,\infty)\to M$ such that $\widetilde{\gamma}_{n}$ limits to $\gamma$ uniformly on compact subsets of $0,\infty)$. Moreover, $\gamma$ is an $S$-maximizer since each $\gamma_{n}$ is an $S$-maximizer. (This follows by upper semicontinuity of the Lorentzian length functional.) Therefore, after a reparameterization, $\gamma$ is a causal geodesic and, in fact, timelike since $S$ is acausal. By future timelike completeness, there is a $c>0$ such that $\gamma(c)\in S_{\tau}$. Set
$q=\gamma(c),\quad q^{\prime}=\gamma(c+1),\quad q_{n}=\widetilde{\gamma}_{n}(c),\quad q_{n}^{\prime}=\widetilde{\gamma}_{n}(c+1).$
Since $b_{n}\to\infty$, we can choose $n$ large enough so that $c+1<b_{n}$. For such large $n$, we have $q_{n},q_{n}^{\prime}\in I^{-}(S_{\tau})$, but $q_{n}^{\prime}\to q^{\prime}\in I^{+}(S_{\tau})$ – yielding an acausality violation of $S_{\tau}$. Hence, $S_{\tau}$ is compact.
If $S_{\tau}$ is not a Cauchy surface, then its Cauchy horizon is nonempty. Suppose, say, $x\in H^{+}(S_{\tau})$. Then, since $S_{\tau}$ is closed and acausal, we can apply [27, Prop. 14.53] to conclude that $x$ is the future endpoint of a past-inextendible null geodesic $\eta\subset H^{+}(S_{\tau})$, but this leads to a strong causality violation,
$\eta\subset J^{-}(x)\cap H^{+}(S_{\tau})\subset J^{-}(x)\cap J^{+}(S_{\tau}),$
since the latter set is compact.
@@
Lastly, we show that $S_{\tau}$ has mean curvature $H\leq\tfrac{n}{\tau}$ in the support sense. Let $p,q,\gamma$ be as in the first paragraph of this proof. In particular, each segment of $\gamma$ is maximal and so there are no cut points to $\gamma(t)$ along $\gamma|_{[t,\tau]}$, for $t>0$. (Recall that a cut point to a timelike geodesic segment corresponds to the first point in which the geodesic is no longer maximizing from its starting point. See chapter 9 of [6]. This ensures the smoothness of the exponential map near the geodesic, before the cut point, from its starting point. See [6, Thm. 9.12] and [27, Prop. 10.10].) Consequently, each future Lorentzian sphere $\Sigma_{r}=\{x\in M\mid d\big(\gamma(t),x\big)=r\}$, for $r\in(0,\tau-t]$, is smooth near $\gamma(r+t)$. It follows that $\Sigma_{\tau-t}$ lies locally to the future of $S_{\tau}$ near $q$, for otherwise there would be a causal curve from $S$ to $S_{\tau}$ with length greater than $\tau$. Hence $\Sigma_{\tau-t}$ is a future support hypersurface for $S_{\tau}$ at $q$. Let $H(r)$ be the mean curvature of $\Sigma_{r}$ at $\gamma(r+t)$. The Raychaudhuri equation [22] (for a hypersurface orthogonal geodesic vector field) gives, when restricted to $\gamma$,
$\frac{dH}{dr}=-{\rm Ric}(\gamma^{\prime},\gamma^{\prime})-\frac{H^{2}}{n}-2\sigma^{2}\leq-\frac{H^{2}}{n}.$ (3.9)
@@
Standard ODE comparison results yield $H(r)\leq\frac{n}{r}$ (see e.g., [24, Section 1.6]), [13, Section 2]). Evaluating at $r=\tau-t$ and letting $t\to 0$ gives the result. ∎
@@
Remark. In fact for $\tau$ small, and provided that $S$ is smooth and spacelike, $S_{\tau}$ will be a Cauchy surface even without the completeness assumption. More concretely, define the “future existence time” $T_{0}>0$ from $S$ to be the largest value (possibly $\infty$) such that $\exp_{p}(tN)$ is defined for all $p\in S$ and all $t<T_{0}$. (Here $N$ is the future timelike unit normal field to $S$.) Then, by essentially the same argument, $S_{\tau}$ will be a Cauchy surface for all $\tau<T_{0}$. In this case, for such an $S$, the conclusion of Proposition 4 holds for all $\tau<T_{0}$.
@@
Let $(M,g)$ be a spacetime with compact Cauchy surfaces satisfying the strong energy condition. Suppose that $S_{1}$ is a Cauchy surface with mean curvature $H_{1}\geq a$ in the support sense and that $S_{2}\subset I^{+}(S_{1})$ is a Cauchy surface with mean curvature $H_{2}\leq b$ in the support sense, for some constants $b<c<a$. Then there exists a smooth spacelike Cauchy surface $S$ between [^5] I.e., $S\subset I^{+}(S_{1})\cap I^{-}(S_{2})$ $S_{1}$ and $S_{2}$ with constant mean curvature $H=c$, which is obtained as the asymptotic limit of mean curvature flow.
@@
Let $V_{0}=F_{0}(V)$ be any embedded smooth spacelike Cauchy surface between $S_{1}$ and $S_{2}$. Consider the evolution problem (2.1). We show that the corresponding flow $V_{s}$, given by (2.2), always lies within the compact region between $S_{1}$ and $S_{2}$. Indeed, suppose there is a point $(x_{0},s_{0})\in V\times[0,\infty)$ where the flow $V_{s}$ touches $S_{2}$ for the first time. At this point $p_{0}:=F(x_{0},s_{0})$, we must have $\tfrac{d}{ds}u\geq 0$, as otherwise the flow $V_{s}$ would necessarily have touched $S_{2}$ at some value $s_{1}<s_{0}$. We conclude from (2.8) that $H(x_{0},s_{0})\geq c$. Since $S_{2}$ has mean curvature $H_{2}\leq b<c$ in the support sense, for each $\varepsilon>0$, there exists a future support hypersurface $W$ for $S_{2}$ at $p_{0}$ with mean curvature $H_{W}\leq b+\varepsilon$. By choosing $\varepsilon$ sufficiently small, it is easy to see that there is a $d\in(b,c)$ such that $H_{V_{s_{0}}}>d$ at $p_{0}$ and $H_{W}<d$ at $p_{0}$. Since $W$ is locally to the future of $S_{2}$ near $p_{0}$ and hence also locally to the future of $V_{s_{0}}$ near $p_{0}$, we arrive at a contradiction to the maximum principle. Thus the flow always remains below $S_{2}$. Likewise, the flow always remains above $S_{1}$.
@@
Since the Cauchy surfaces are compact, we can find $t$-slices of the temporal function 2.5 such that $S_{1}$ always lies above a slice $t=\alpha$ and $S_{2}$ lies below a slice $t=\beta$. Therefore the flow always remains in the smooth compact region of spacetime between $t=\alpha$ and $t=\beta$. It follows via [12, Thms. 4.1 and 4.2] that the flow $V_{s}$ is defined for all $s\in[0,\infty)$ and (by taking a suitable subsquence) limits to a smooth spacelike Cauchy surface $S$ with constant mean curvature $H=c$. ∎
@@
Remark. As alluded to in the introduction, Proposition 5 extends Theorem 4.2 in [12] to the case of mean curvature barriers in the support sense. An existence result similar to [12, Thm. 4.2] was obtained by Claus Gerhardt [20], again assuming smooth barriers. Gerhardt’s proof was via a variational method and did not require the Ricci curvature condition. There appears to be certain technical issues in trying to extend Gerhardt’s proof to the case of barriers in the support sense. A closely related situation arises in [1]. We thank the authors of [1] for a discussion pertaining to some of these issues.
@@
Proof of Theorem 3. By assumption, there is a smooth spacelike Cauchy surface $V$ with mean curvature $H_{V}\geq 0$. If $H_{V}=0$ identically, then we are done. Otherwise, there is a point $p\in V$ with $H_{V}(p)>0$. We can perturb $V$ slightly to the past to produce a smooth spacelike Cauchy surface with strictly positive mean curvature. To see this, let $V_{r}$ be a variation of $V_{0}=V$ with variation vector $\phi\nu$, where $\phi$ is a smooth function on $V$ and $\nu$ is the future directed unit timelike normal to $V$. Let $H_{r}$ be the mean curvature of $V_{r}$. A standard computation shows that
@@
$\frac{\partial H_{r}}{\partial r}\bigg|_{r=0}\,=\,-L\phi\quad\text{ where }\quad L\phi\,:=\,-\Delta\phi+\big({\rm Ric}(\nu,\nu)+\tfrac{H_{0}^{2}}{n}+\sigma^{2}\big)\phi,$
@@
where $\Delta$ is the spatial Laplacian on $V$ and $\sigma^{2}$ is the norm squared of the trace-free part of the second fundamental form of $V$. Since $H_{0}(p)>0$, the Rayleigh quotient formula for the principal eigenvalue $\lambda_{1}$ of $L$,
[…]
shows that $\lambda_{1}>0$. Then, by varying with respect to the principal eigenfunction, there are times $r<0$ such that $H_{r}>0$ everywhere on $V_{r}$.
@@
Thus we can assume – without loss of generality – that $H_{V}>0$. Since $V$ is smooth, we can choose constants $a>c>0$ such that $H_{V}\geq a$. By taking $\tau$ sufficiently large, Proposition 4 implies the existence of a Cauchy surface $V_{\tau}$ with mean curvature $H_{\tau}\leq b$ in the support sense, for some $b<c$. Then Proposition 5 implies that there is a smooth spacelike Cauchy surface between $V$ and $V_{\tau}$ with constant mean curvature $H=c$. ∎
@@
Remark. In Theorem 3, we showed the existence of a smooth Cauchy surface, say $V_{1}$, with constant mean curvature $H=c_{1}>0$ (in the case that $V$ is not maximal); moreover, $V_{1}$ is unique since $c_{1}\neq 0$, see e.g., [20, Thm. 7.3]. In fact, we can say more. By another application of Theorem 3, for any $0<c_{2}<c_{1}$, we can find a CMC Cauchy surface $V_{2}\subset I^{+}(V_{1})$ with mean curvature $H=c_{2}$. Then the region between $V_{1}$ and $V_{2}$ is covered by unique CMC Cauchy surfaces with mean curvatures $c\in(c_{2},c_{1})$, as follows from [20, Thm. 7.5].